OpenStudy (anonymous):

Graphf(x)=2x^2 – 4. And find the inverse

OpenStudy (anonymous):

Plot some x values to f(x).

OpenStudy (anonymous):

Example: f(x)=2x&2-4 f(0)=2(0)^2-4; so y=-4